# A Non-Parametric Sensitivity Analysis for Bounding Bias in Hybrid Control Trials

### v1

@@ Abstract
In the digital era it is easier than ever to collect and exploit rich covariate information in trials. Recent work explores how to use this information to integrate external controls, including the use of hybrid control trials (HCTs) where a randomized controlled trial is augmented with external controls. HCTs are of particular interest due to their ability to preserve partial randomization while also improving trial efficiency. However, most HCT estimators rely on an unrealistic assumption: that the external controls are drawn from the same population as the trial subjects (perhaps conditionally on covariates). There has been little formal work to quantify the inevitable bias introduced from a violation of this assumption, slowing the acceptance of HCT designs. To address this, we introduce a non-parametric sensitivity analysis that recognizes that the assumption can be reframed as a ‘no unobserved confounders’ assumption. We leverage omitted variable bias methodologies to estimate the maximum bias introduced from unmeasured covariates, allowing for a critical evaluation of the causal gap that invalidates significant findings. We show that with a relatively weak understanding of the covariate-outcome relationship and of the distinguishability of trial and external subjects, this method reliably bounds bias while also allowing for gains in efficiency. We conclude by discussing considerations for designing and evaluating HCTs, drawing on insights from simulations and theoretical analyses.
@@ 1 Introduction
Randomized controlled trials (RCTs) are widely regarded as the gold standard in causal inference because they eliminate confounding, allowing for unbiased estimation [37]. However, RCTs face many limitations, including high costs, ethical constraints, and recruitment difficulties, that make them infeasible or impossible at times [10, 19, 39]. These challenges can compromise research quality through underpowering of trials or inflated type I error rates. Such problems are particularly common in research areas such as rare diseases or vulnerable populations, including pregnant mothers and marginalized populations [6, 23].
@@
Hybrid control trials (HCTs) are an alternative to randomized controlled trials in which an RCT is supplemented with additional control data from external sources (also known as “pooling”). Although not common in practice, the use of HCTs has been proposed as early as 1976 but has since received attention due to their potential to alleviate some of the restrictions mentioned above on RCTs [34, 38, 9]. Regulatory agencies, including the FDA and EMA, have shown a growing interest and commitment to real-world evidence from these trials [6, 34, 9, 1, 4, 5, 3]. The 21st Century Cures Act likewise acknowledges the need to accelerate medical product development in the face of lengthy randomized controlled trials [2]. ‘Borrowing’ information from external sources aligns with these goals and can decrease trial timelines and sample sizes. Furthermore, data availability continues to grow as trials are conducted and pharmaceutical companies participate in data sharing of concurrent control data. Using external controls also offers the opportunity to change the randomization ratio in the RCT portion of the HCT [30]. This allows for more patients to receive the treatment rather than the control, making the trial both more ethical and attractive to prospective participants. With the growth of data, increased support from regulatory agencies, and the existing challenges of traditional RCTs, the time has come to explore data borrowing alternatives such as hybrid control trials.
@@
Multiple data borrowing methods have been proposed with different positions on the bias-variance tradeoff spectrum. At one end of the spectrum, non-HCT methods such as prognostic covariate adjustment train models on previous data to create prognostic scores for trial participants [36, 29]. These do not directly incorporate external data as additional control data but instead use the predictions as an additional covariate to adjust on. Thus, bias is not introduced with the use of the external data. However, since information is borrowed conservatively, the variance reductions are relatively small compared to other methods.
Some HCT methods incorporate external data more directly by ‘pooling’ external and trial control data together. By treating external data as trial data, researchers artificially increase the RCT sample size. Typically, variance decreases and power increases with an increase in sample size. However, the risk of bias is also substantially increased if and when the external population does not truly reflect the trial population. These methods are often referred to as “static” borrowing: gains in efficiency are highest, but robustness against bias is lowest. With static borrowing, researchers always pool the data, assuming that the external data is compatible [42]. The artificial trial sample size increases accordingly to the size of the external sample size, rapidly improving trial power. However, assuming compatibility of the data (which is often unrealistic) introduces estimation bias and type I error inflation.
@@
Even more drastic borrowing methods exist such as single-arm trials where trial participants receive only the treatment, and the control arm consists solely of external data. Note that this is not a form of an HCT as the control arm is not hybrid. Single-arm trials are only acceptable in rare cases since they rely fully upon external data to make comparisons of treatment against control, making them the most susceptible to bias from the use of external data [3, 7].
@@
Besides these “static” borrowing approaches, researchers have also considered “dynamic” approaches [38]. Dynamic borrowing methods fall between the extremes of the bias-variance spectrum, assessing compatibility of control data to adaptively change reliance on external data. These include Bayesian approaches such as power priors and the commensurate prior model [38, 16, 33, 20]. Bayesian power prior methods rely on a weighting power parameter that ranges from 0 (equivalent to an RCT with no external data) to 1 (equivalent to a fully pooled HCT where external and trial data are treated equally). This power parameter is given a prior distribution that then is updated based on the utilized data to optimize weighting of data. These approaches are attractive since they measure differences in the data while still cleverly incorporating external control data. However, large power gains generally cannot be realized without sacrificing type I error control [24, 18].
@@
Frequentist dynamic borrowing includes the test-then-pool approach where data is first assessed for significant differences such as by hypothesis tests [38]. If differences are significant, no pooling occurs; otherwise, pooling proceeds. By testing beforehand, researchers attempt to minimize bias by only pooling when data incompatibility is statistically insignificant. However, if no pooling occurs, researchers solely rely on RCT data with no avenues for gains in efficiency. If pooling occurs, the data are treated as if they are additional trial control units, which results in considerable power gains. However, this approach is criticized for its inability to consistently detect biases, leaving it vulnerable to bias from pooling [27]. Finally, more rigorous frequentist dynamic borrowing methods have been proposed that rely on adaptive lasso penalization to select a subset of suitable external controls [17]. Only external controls whose bias parameters are (or close to) 0 are utilized, resulting in minimized bias from pooling. This method works well compared to others, notably in terms of mean-squared error, especially when differences in control groups are large. However, issues with type I error control arise with finite samples, and gains in power are not as large as other pooling methods due to the conservative nature of selecting suitable controls when differences are already small.
@@
Due to the large potential to increase power, we focus on HCT pooling methods that treat external data similarly to trial data (these include static and frequentist dynamic borrowing methods). The trial-specific average treatment effect (ATE) is the primary parameter of interest in these pooled data experiments, and thus many estimators have been proposed in HCT literature [41, 28, 25]. However, these HCT estimators rely on an unrealistic assumption: mean exchangeability of the controls [28]. A violation of this assumption introduces bias into estimation, potentially inflating type I error rates or decreasing power depending on the direction of the bias. In the context of clinical trials — and particularly of phase III confirmatory trials — these errors are especially problematic. Literature has focused on the development of these estimators, but few, if any, have discussed a formal approach to quantify the bias introduced by an inevitable violation of this assumption, slowing the acceptance of HCTs [41, 1].
@@ 1.1 Contributions
To mitigate against this bias, we propose that stringent sensitivity analysis has the potential to maintain the reliability of findings from these trials. We extend research from Chernozhukov et al. [14] where a non-parametric sensitivity analysis leverages efficient debiased machine learning and produces easily interpretable bounds on omitted variable bias in observational studies. In theory, the mean exchangeability assumption is met when all relevant covariates are controlled for confounding. It follows that a violation of the mean exchangeability assumption is a result of omitted variable bias. We modify this sensitivity analysis in the context of HCTs to reliably bound bias given plausible increases in the explanatory power from the omitted variables on the “Riesz representer” and outcome regression of the trial-specific ATE without making any additional assumptions on the data generating process [14]. Results of the sensitivity analysis can then be compared against the strength of evidence from HCTs to determine whether conclusions are still significant even after accounting for the potential bias.
@@
In Section 2 we begin by defining necessary notation and preliminaries for the hybrid control trial setting. We also explore a doubly robust efficient estimator for the trial-specific average treatment effect [28]. In Section 3 we validate that it is possible to extend the omitted variable bias sensitivity analysis to the HCT setting. Through identification of the Riesz representer and outcome regression of the trial-specific ATE, we derive an estimator for the bias bound $B$. In Section 4, we use simulations to demonstrate that the risk of bias and error is minimized with the sensitivity analysis and that trial efficiency can be improved even when bias is present. We do this by showing (i.) that the true bias is reliably bounded by $B$, (ii.) that the sensitivity analysis corrects inflated type I error rates in HCTs, and (iii.) that power can be increased when using the HCT estimator combined with the sensitivity analysis. In Section 5, we summarize key findings and explore guidelines for designing and evaluating HCTs based on these results as well as the potential impacts of this research.
@@ 2.1 Notation
Consider a hypothetical hybrid control trial, consisting of a randomized controlled trial and an external control arm. Let the variable $D$ denote trial participation with $D=1$ indicating that a subject was a participant in the randomized controlled trial and $D=0$ indicating that a subject was not a participant in the randomized controlled trial. Equivalently, $D=0$ implies the participant is part of the external control arm. Let the variable $Y$ represent the outcome of interest in the study and $X$ represent the measured covariate(s) in the study. Let $A$ represent the binary treatment of interest with $A=1$ indicating that a subject received the treatment of interest and $A=0$ indicating that a subject did not receive the treatment. Note that since the external data only consist of control data, a subject with $D=0$ implies $A=0$. The observed data follow structure $O=(D,A,X,Y)$. Furthermore, we assume that the data $O_{1},...,O_{n}$ are i.i.d. distributed from a true data distributing distribution $\operatorname*{\mathbb{I\kern-1.49994ptP}}_{0}$.
@@
$A$ Treatment indicator: $A=1$ if treated, $A=0$ otherwise.
@@
$Z$ Unobserved confounders (s.t. $Y_{a}\perp\!\!\perp A\mid X,Z$).
@@
We also consider the unobserved full dataset. Let $Y_{1}$ and $Y_{0}$ denote the counterfactual outcomes where $Y_{a}$ represents the outcome that would be observed under treatment $a\in\{0,1\}$. We define the observed outcome $Y$ under treatment $A$ with the counterfactual outcomes $Y_{a}$ such that $Y=A\cdot Y_{1}+(1-A)\cdot Y_{0}$ (others call this the “causal consistency” assumption, often bundled with the IID assumption as “SUTVA”). Assume that there is a set of unobserved random variables Z such that $Y_{a}\perp\kern-9.0pt\perp A|X,Z$. These represent our “unobserved confounders” without which we cannot guarantee mean exchangeability. The full data structure follows as $O^{*}=(D,A,X,Z,Y_{1},Y_{0})$, which is i.i.d. distributed from some distribution $\operatorname*{\mathbb{I\kern-1.49994ptP}}_{0}^{*}$. We can think of $\operatorname*{\mathbb{I\kern-1.49994ptP}}_{0}$ as marginal relative to $\operatorname*{\mathbb{I\kern-1.49994ptP}}_{0}^{*}$ over $Z$ and with only $Y$ observed rather than $Y_{1}$ and $Y_{0}$.
Let $N_{d}$ with $d\in\{0,1\}$ denote the sample size of the trial and external datasets for $d=1$ and $d=0$ respectively. Since we are assuming $D$ is i.i.d. distributed from $\operatorname*{\mathbb{I\kern-1.49994ptP}}_{0}$, the external sample size and RCT sample sizes are not fixed. The total sample size of the hybrid control trial $n$ is the sum of $N_{0}$ and $N_{1}$.
@@
$\mu(A,X)$ $E\bigl{[}Y\mid X,D=1,A=a\bigr{]}$, the trial‐specific outcome regression.
@@ 2.2 The Trial-Specific Average Treatment Effect
In this paper, we focus on the trial-specific average treatment effect. This measures the causal effect of the treatment $A$ on outcomes $Y$ for the population of people who were eligible to participate in the randomized controlled trial. Although other estimands of interest exist for hybrid control trials, we focus on the trial-specific average treatment effect due to its prevalence in hybrid control trial causal inference literature [28, 41, 40]. This parameter is also of key importance in the pharmaceutical and regulatory context where RCT trial recruitment is based on a carefully selected population, so inference would likewise need to be focused on this specific population.
@@
Strong ignorability is a common assumption to identify the average treatment effect when conducting inference in observational studies. That condition is met by design for the purposes of our paper since the trial is randomized. On the other hand, the mean exchangeability assumption is not addressed by randomization, leaving a gap between the observable estimand that can be estimated and the causal estimand of scientific interest. This gap is precisely the bias induced by pooling in HCTs.

### v2

@@ Abstract
We study hybrid control trials (HCTs), in which a randomized controlled trial (RCT) is augmented with external control patients. Existing approaches for HCTs typically assume conditional exchangeability of the concurrent and external controls to identify trial-specific effects. When violated, this can induce substantial unquantified bias, which in turn limits the acceptability of HCTs in regulatory settings. We treat violations of mean exchangeability as omitted variable bias and develop a non-parametric sensitivity analysis that (i) applies to the efficient, doubly robust HCT estimator of the trial-specific ATE, and (ii) delivers sharp bounds on the bias induced by unmeasured covariates. Building on recent work in double machine learning, our approach characterizes the maximal bias in terms of two partial R-squared sensitivity parameters: the additional explanatory power that unmeasured confounders could have for the outcome regression and for trial participation. For any given choice of these parameters, we construct valid confidence bounds for bias-adjusted treatment effects and visualize critical causal gaps via contour plots and robustness values that show how strong unmeasured confounding would need to be to overturn nominally significant HCT findings. Through simulations, we show that the method (i) reliably upper-bounds true bias, (ii) restores type I error control in settings where naïve HCT analysis is anti-conservative, and (iii) can still deliver meaningful power gains and RCT sample-size reductions even under moderate violations of mean exchangeability. We illustrate the approach in a phase III trial on diabetes, supplemented with external controls. We discuss practical guidelines for designing and evaluating HCTs, including external-data selection, sample-size allocation, and interpretation of sensitivity contours.
@@ 1 Introduction
Randomized controlled trials (RCTs) are widely regarded as the gold standard in causal inference because they eliminate confounding, allowing for unbiased estimation [sox_methods_2012]. However, RCTs face many limitations, including high costs, ethical constraints, and recruitment difficulties, that make them infeasible or impossible at times [bentley_conducting_2019, glennerster_chapter_2017, temple_placebo-controlled_2000]. These challenges can compromise research quality through underpowering of trials. Such problems are particularly common in research areas such as rare diseases or vulnerable populations, including pregnant mothers and marginalized populations, as well as in Phase II trials [noauthor_rare_2019, jahanshahi_use_2021].
@@
Hybrid control trials (HCTs) are an alternative to randomized controlled trials in which an RCT is supplemented with additional control data from external sources (also known as “pooling”). Although not common in practice, the use of HCTs has been proposed as early as 1976 but has since received attention due to their potential to alleviate some of the restrictions mentioned above on RCTs [pocock_combination_1976, tan_augmenting_2022, baumfeld_andre_trial_2020]. Regulatory agencies, including the FDA and EMA, have shown a growing interest and commitment to real-world evidence from these trials [noauthor_rare_2019, pocock_combination_1976, baumfeld_andre_trial_2020, noauthor_international_2000, noauthor_fda_2019, noauthor_fda_2019-2, noauthor_framework_2018]. The 21st Century Cures Act likewise acknowledges the need to accelerate medical product development in the face of lengthy randomized controlled trials [noauthor_21st_2017]. ‘Borrowing’ information from external sources aligns with these goals and can decrease trial timelines and sample sizes. Furthermore, data availability continues to grow as trials are conducted and pharmaceutical companies participate in data sharing of concurrent control data. Using external controls also offers the opportunity to change the randomization ratio in the RCT portion of the HCT [lim_minimizing_2018]. This allows for more patients to receive the treatment rather than the control, making the trial both more ethical and attractive to prospective participants. With the growth of data, increased support from regulatory agencies, and the existing challenges of traditional RCTs, the time has come to explore data borrowing alternatives such as hybrid control trials.
@@
Multiple data borrowing methods have been proposed with different positions on the bias-variance tradeoff spectrum. At one end of the spectrum, non-HCT methods such as prognostic covariate adjustment train models on previous data to create prognostic scores for trial participants [schuler_increasing_2022, liao_prognostic_2025]. These do not directly incorporate external data as additional control data but instead use the predictions as an additional covariate to adjust on. Thus, bias is not introduced with the use of the external data. However, since information is borrowed conservatively, the variance reductions are relatively small compared to other methods.
@@
At the other end of the spectrum, the most drastic borrowing method is the single arm trial where trial participants receive only the treatment, and the control arm consists solely of external data. Note that this is not a form of an HCT as the control arm is not hybrid. Single-arm trials are only acceptable in rare cases since they rely fully upon external data to make comparisons of treatment against control, making them the most susceptible to bias from the use of external data [noauthor_framework_2018, agrawal_use_2023].
HCT methodologies lie in the middle of the spectrum with their positions determined by how they incorporate the external data. Some HCT methods incorporate external data more directly by ‘pooling’ external and trial control data together. By treating external data as trial data, researchers artificially increase the RCT sample size. Typically, variance decreases and power increases with an increase in sample size. However, the risk of bias is also substantially increased if and when the external population does not truly reflect the trial population. These methods are often referred to as “static” borrowing: gains in efficiency are highest, but robustness against bias is lowest. With static borrowing, researchers always pool the data, assuming that the external data is compatible [viele_use_2014]. The artificial trial sample size increases accordingly to the size of the external sample size, rapidly improving trial power. However, assuming compatibility of the data (which is often unrealistic) introduces estimation bias and type I error inflation.
@@
Besides these “static” borrowing approaches, researchers have also considered “dynamic” approaches [tan_augmenting_2022]. Dynamic borrowing methods fall between the extremes of the bias-variance spectrum, assessing compatibility of control data to adaptively change reliance on external data. These include Bayesian approaches such as power priors and the commensurate prior model [tan_augmenting_2022, duan_evaluating_2006, neuenschwander_note_2009, hobbs_hierarchical_2011]. Bayesian power prior methods rely on a weighting power parameter that ranges from 0 (equivalent to an RCT with no external data) to 1 (equivalent to a fully pooled HCT where external and trial data are treated equally). This power parameter is given a prior distribution that then is updated based on the utilized data to optimize weighting of data. These approaches are attractive since they measure differences in the data while still cleverly incorporating external control data. However, large power gains generally cannot be realized without sacrificing type I error control [koppschneider_power_2020, ghadessi_roadmap_2020].
@@
Frequentist dynamic borrowing includes the test-then-pool approach where data is first assessed for significant differences such as by hypothesis tests [tan_augmenting_2022]. If differences are significant, no pooling occurs; otherwise, pooling proceeds. By testing beforehand, researchers attempt to minimize bias by only pooling when data incompatibility is statistically insignificant. However, if no pooling occurs, researchers solely rely on RCT data with no avenues for gains in efficiency. If pooling occurs, the data are treated as if they are additional trial control units, though still subject to some weighting imposed by the estimator. This can produce substantial power gains. However, this approach is criticized for its inability to consistently detect biases, leaving it vulnerable to bias from pooling [li_revisit_2020]. Finally, more rigorous frequentist dynamic borrowing methods have been proposed that rely on adaptive lasso penalization to select a subset of suitable external controls [gao_improving_2024]. Only external controls whose bias parameters are (or close to) 0 are utilized, resulting in minimized bias from pooling. This method works well compared to others, notably in terms of mean-squared error, especially when differences in control groups are large. However, issues with type I error control arise with finite samples, and gains in power are not as large as other pooling methods due to the conservative nature of selecting suitable controls when differences are already small.
@@
Due to the large potential to increase power, we focus on a method that allows for pooled control data. The trial-specific average treatment effect (ATE) is the primary parameter of interest in many data-borrowing experiments, and thus many estimators have been proposed in HCT literature [valancius_causal_2024, li_improving_2023, laan_targeted_2011]. However, these HCT estimators rely on an unrealistic assumption: mean exchangeability of the controls [li_improving_2023]. A violation of this assumption introduces bias into estimation, potentially inflating type I error rates or decreasing power depending on the direction of the bias. In the context of clinical trials — and particularly of phase III confirmatory trials — these errors are especially problematic. Literature has focused on the development of these estimators, but few, if any, have discussed a formal approach to quantify the bias introduced by an inevitable violation of this assumption, slowing the acceptance of HCTs [valancius_causal_2024, noauthor_international_2000].
@@ 1.1 Contributions
To mitigate against this bias, we propose that stringent sensitivity analysis has the potential to maintain the reliability of findings from these trials. We extend research from chernozhukov_long_2024 where a non-parametric sensitivity analysis leverages efficient debiased machine learning and produces easily interpretable bounds on omitted variable bias in observational studies. In theory, the mean exchangeability assumption is met when all relevant covariates are controlled for confounding. It follows that a violation of the mean exchangeability assumption is a result of omitted variable bias. We modify this sensitivity analysis in the context of HCTs to reliably bound bias given plausible increases in the explanatory power from the omitted variables on the “Riesz representer” and outcome regression of the trial-specific ATE [chernozhukov_long_2024]. In doing so, we examine robustness of results against user-supplied sensitivity parameters, which can be evaluated for plausibility with observed variables, instead of relying on the untestable mean exchangeability assumption for inference. Results of the sensitivity analysis can then be compared against the strength of evidence from HCTs to determine whether conclusions are still significant even after accounting for the potential bias.
@@
In Section 2 we begin by defining necessary notation and preliminaries for the hybrid control trial setting. We also explore a doubly robust efficient estimator for the trial-specific average treatment effect [li_improving_2023]. In Section 3 we validate that it is possible to extend the omitted variable bias sensitivity analysis to the HCT setting. Through identification of the Riesz representer and outcome regression of the trial-specific ATE, we derive an estimator for the bias bound $B$. In Section 4, we use simulations to demonstrate that the risk of bias and error is minimized with the sensitivity analysis and that trial efficiency can be improved even when bias is present. We do this by showing (i.) that the true bias is reliably bounded by $B$, (ii.) that the sensitivity analysis corrects inflated type I error rates in HCTs, and (iii.) that power can be increased when using the HCT estimator combined with the sensitivity analysis. In Section 5, we provide a case study using data from prior clinical trials conducted by Novo Nordisk (AS). We demonstrate how to apply the sensitivity analysis to results from an HCT study and discuss important practical considerations. In Section 6, we summarize key findings and explore guidelines for designing and evaluating HCTs based on these results as well as the potential impacts of this research.
@@ 2.1 Notation
Consider a hypothetical hybrid control trial, consisting of a randomized controlled trial and an external control arm. Let the variable $D$ denote trial participation with $D=1$ indicating that a subject was a participant in the randomized controlled trial and $D=0$ indicating that a subject was not a participant in the randomized controlled trial. Equivalently, $D=0$ implies the participant is part of the external control arm. Let the variable $Y$ represent the outcome of interest in the study and $X$ represent the measured covariate(s) in the study. Let $A$ represent the binary treatment of interest with $A=1$ indicating that a subject received the treatment of interest and $A=0$ indicating that a subject did not receive the treatment. Note that since the external data only consist of control data, a subject with $D=0$ implies $A=0$. The observed data follow structure $O=(D,A,X,Y)$. Furthermore, we assume that the data $O_{1},...,O_{n}$ are i.i.d. distributed from a true data distributing distribution $\operatorname*{\mathbb{P}}_{0}$.
@@
AA Treatment indicator: $A=1$ if treated, $A=0$ otherwise.
@@
$Z$ Unobserved confounders (s.t. $Y_{0}\perp\!\!\perp D\mid X,Z$).
@@
We also consider the unobserved full dataset. Let $Y_{1}$ and $Y_{0}$ denote the counterfactual outcomes where $Y_{a}$ represents the outcome that would be observed under treatment $a\in\{0,1\}$. We define the observed outcome $Y$ under treatment $A$ with the counterfactual outcomes $Y_{a}$ such that $Y=A\cdot Y_{1}+(1-A)\cdot Y_{0}$ (others call this the “causal consistency” assumption). Assume that there is a set of unobserved random variables Z such that $Y_{0}\perp\kern-9.0pt\perp D|X,Z$. These represent our “unobserved confounders” without which we cannot guarantee mean exchangeability. The full data structure follows as $O^{*}=(D,A,X,Z,Y_{1},Y_{0})$, which is i.i.d. distributed from some distribution $\operatorname*{\mathbb{P}}_{0}^{*}$. We can think of $\operatorname*{\mathbb{P}}_{0}$ as marginal relative to $\operatorname*{\mathbb{P}}_{0}^{*}$ over $Z$ and with only $Y$ observed rather than $Y_{1}$ and $Y_{0}$.
Let $N_{d}$ with $d\in\{0,1\}$ denote the sample size of the trial and external datasets for $d=1$ and $d=0$ respectively. Since we are assuming $D$ is i.i.d. distributed from $\operatorname*{\mathbb{P}}_{0}$, the external sample size and RCT sample sizes are not fixed. The total sample size of the hybrid control trial $n$ is the sum of $N_{0}$ and $N_{1}$.
@@
$\mu(A,X)$ $E\bigl[Y\mid X,D=1,A=a\bigr]$, the trial‐specific outcome regressiona.
@@
  • a
    Under mean exchangeability discussed in the following section, $\mu(0,X)=E[Y\mid X,A=0]$.
@@ 2.2 The Trial-Specific Average Treatment Effect
In this paper, we focus on the trial-specific average treatment effect. This measures the causal effect of the treatment $A$ on outcomes $Y$ for the population of people who were eligible to participate in the randomized controlled trial. Although other estimands of interest exist for hybrid control trials, we focus on the trial-specific average treatment effect due to its prevalence in hybrid control trial causal inference literature [li_improving_2023, valancius_causal_2024, ung_combining_2024]. This parameter is also of key importance in the pharmaceutical and regulatory context where RCT trial recruitment is based on a carefully selected population, so inference would likewise need to be focused on this specific population.
@@
Strong ignorability is a common assumption to identify the average treatment effect when conducting inference in observational studies. That condition is met by design for the purposes of our paper since the trial is randomized. On the other hand, the mean exchangeability assumption is not addressed by randomization, leaving a gap between the observable estimand that can be estimated and the causal estimand of scientific interest. This gap is precisely the bias induced by pooling in HCTs. For the time being we will proceed as if we believed this assumption. Later we will address its impact.